# Neutrino masses from new Weinberg-like operators: Phenomenology of TeV scalar multiplets

### v1

@@ 3.2.1 The $\mathbf{A_{I}}$ scenario
As expected, we obtain a pure massless would-be Nambu-Goldstone boson $z$, associated with the spontaneous breaking of the hypercharge, and a massive pseudo-Nambu-Goldstone boson $J$, related to the explicit breaking of the lepton number in one unit, that we identify as the pseudo-Majoron. Notice that its mass term depends on the LNV quartic couplings of the potential and, as already mentioned, in the limit $(\lambda_{6},\lambda_{7},\lambda_{8})\rightarrow 0$, it becomes massless. The mixing angle $\alpha$ is given by
@@ 3.2.2 The $\mathbf{B_{I}}$ scenario
We now study the case of the $\mathbf{B_{I}}$ scenario as an example. We choose this specific class-B scenario since it is a natural extension of the $\mathbf{A_{I}}$ scenario discuss above. Indeed, in $\mathbf{B_{I}}$ we add to $\mathbf{A_{I}}$ a second multiplet, $\Phi_{2}=\mathbf{4}_{-3/2}$. Moreover, the study of such a scenario is particularly interesting since we can infer the results on the Majoron of the scenario $\mathbf{A_{II}}$ setting to zero all the scalar potential couplings in which $\Phi_{1}$ appears. In this scenario, as in all class-B scenarios, there are three $U(1)$ symmetries: $U(1)_{Y}$ of hypercharge, $U(1)_{L}$ of lepton number, and an additional $U(1)_{X}$ accidental symmetry. In the limit in which all the violating terms in the potential go to zero, we expect to have three massless Nambu-Goldstone bosons; if instead we turn on the $\lambda_{1}$ (Eq. (16)) we get a massive Nambu-Goldstone boson i.e., the pseudo-Majoron $J$, associated with the explicit breaking of the lepton number. If the $\lambda_{i}$ for $i=2,...,9$ are also different from zero (see Table 3), a second massive pseudo-Nambu-Goldstone boson $\Omega$ is obtained, related to the $U(1)_{X}$ accidental explicit symmetry breaking. As in Eq. (20), we can split the neutral scalar fields into their real and imaginary parts:
@@
As expected, we obtain a pure massless would-be Nambu-Goldstone boson $z$, associated with the spontaneous breaking of hypercharge, and two massive pseudo-Nambu-Goldstone bosons associated with the explicit breaking of the $U(1)_{L}$ and $U(1)_{X}$ symmetries, respectively. Even if at the leading order there is no dependence on the LNV potential coupling $\lambda_{1}$, in the limit of $U(1)_{X}$ conservation, namely when $\lambda_{i}$ with $i=2,...,9$ are zero, the pseudo-Majoron $J$ remains massive with $m_{J}^{2}\simeq\lambda_{1}v^{2}(v_{1}^{2}+v_{2}^{2})/(2\,v_{1}^{2}\,v_{2}^{2}\sqrt{3})$. Defining $(\tan\alpha)_{uv}$ as the tangent of mixing between the CP-odd scalars $u$ and $v$, we obtain:
@@
It is interesting to notice that, whenever a massive pseudoscalar is included in our scenario, we might set lower bounds on the LNV couplings. Indeed, such particles cannot have a mass which is below 45 GeV, otherwise we would have observed them in $Z$-boson decays CMS:2022ett . If we have terms proportional to $\lambda{{}^{\prime\prime}}\Phi_{i}H^{3}$ in the potential ($\mathbf{A_{I}}$, $\mathbf{A_{II}}$, $\mathbf{B_{I}}$ scenarios), neglecting small corrections proportional to $v_{\Phi}$, we have
@@
This means that the lower limit on $\lambda^{\prime\prime}$ is rather stringent: $\lambda^{\prime\prime}>10^{-5}$. In other scenarios, where we can have terms like $\lambda_{1}\Phi_{1}\Phi_{2}H^{2}$ and $\lambda^{\prime}\Phi_{i}^{2}H^{2}\,(i=1,2)$, there is no enhancement from $v_{i}$ in the denominator, and the requirement $m_{J}>45$ GeV gives $\lambda_{1},\lambda^{\prime}>10^{-1}$.
@@ 3.3 The VEVs of new scalars and the $\rho$ parameter
Table 4: 95% CL upper limits on the VEVs of the new scalar multiplet from the $\rho$ parameter and the possibility of inducing small VEVs in class-A scenarios.
First, let us consider the scenarios belonging to the class-A, namely the ones in which we add a single scalar multiplet: $\Phi_{1}=\mathbf{4}_{-1/2}$ for $\mathbf{A_{I}}$ and $\Phi_{1}=\mathbf{4}_{-3/2}$ for $\mathbf{A_{II}}$. Recalling the general formula, Eq. (29), we get
@@
Using the experimental bound, we obtain the limits on the new VEVs, which are listed in Table 4. As expected, the BSM VEVs are always much smaller than the SM Higgs one. Using this result for which $v_{1}\sim\mathcal{O}(1\,\mathrm{GeV})$, it is also possible to estimate the energy scale of the UV complete theory. Indeed, since from the Weinberg-like operators the neutrino mass can be written as $m_{\nu}\sim vv_{1}/\Lambda$ and $m_{\nu}\sim v_{1}^{2}/\Lambda$ (if the dimension-5 operators $\mathcal{O}_{5}^{(1)}$ and $\mathcal{O}_{5}^{(2)}$ are generated), imposing $m_{\nu}\gtrsim\mathcal{O}(0.05\,\mathrm{eV})$ we can estimate the order of magnitude of the energy scale $\Lambda$. Since we do not know in principle which of the operators gives the leading contribution to neutrino masses, we show here the mass scales $\Lambda^{(i)}$ one would obtain if only one of the operators exists. Taking the Wilson coefficients to be $\mathcal{O}(1)$, we obtain for perturbative Wilson coefficients $\Lambda^{(1)}\lesssim 10^{12}$ GeV for $\mathcal{O}_{5}^{(1)}$, which is present for both scenarios $\mathbf{A_{I,II}}$ and $\Lambda^{(2)}\lesssim 10^{10}$ GeV for $\mathcal{O}_{5}^{(2)}$, which is present for the scenario $\mathbf{A_{I}}$ but not for the scenario $\mathbf{A_{II}}$.
@@
Figure 4: Allowed regions for the new scalar multiplets VEVs in class-$\mathbf{B}$ scenarios computed using the 95% CL limits on the $\rho$ parameter, $\rho=1.00017\pm 0.00025$ ParticleDataGroup:2022pth . Two different structures emerge: hyperbolae for $\mathbf{B_{I,III,V,VI}}$ and ellipses for $\mathbf{B_{II,IV}}$, see text for details.
@@
When the allowed region is an ellipse (scenarios $\mathbf{B_{II}}$ and $\mathbf{B_{IV}}$), the values of the two semi-axes $a_{1,2}$ correspond to the maximum values that the VEVs can assume. On the other hand, when the allowed region is the space between two hyperbolae (scenarios $\mathbf{B_{I}}$, $\mathbf{B_{III}}$, $\mathbf{B_{V}}$ and $\mathbf{B_{VI}}$), the BSM VEVs can in principle also have large values if there is some partial cancellation among the contributions. However, they cannot exceed the limit imposed by the Fermi constant, namely 174 GeV. Moreover, given that the mass of the top quark is proportional to the SM Higgs VEV $v$, by perturbativity the latter is not allowed to be too small. Thus, the new VEVs $v_{1}$ and $v_{2}$ also for the $\mathbf{B_{I,III,V,VI}}$ scenarios cannot exceed roughly $\mathcal{O}(10\,\text{GeV})$.
Analogous to the case of class-A scenarios, we can deduce the energy scale of the new physics imposing that neutrino masses are $m_{\nu}\gtrsim\mathcal{O}(0.05\,\mathrm{eV})$ and that the Wilson coefficients are of $\mathcal{O}(1)$. Since for class-B scenarios we can have the dimension-5 operator $\mathcal{O}_{5}^{(3)}$ from which $m_{\nu}\sim v_{1}v_{2}/\Lambda$, we obtain for all the scenarios, an energy scale $\Lambda\lesssim\mathcal{O}(10^{10})$ GeV. The semi-axes of the conics which represent the allowed region for the new scalar VEVs obtained from EWPT for the class-B scenarios are listed in Table 5. Notice that in scenario $\mathbf{B_{I}}$ the Wilson coefficients for $\mathcal{O}_{5}^{(1)}$ and $\mathcal{O}_{5}^{(2)}$ related to neutrino masses are not zero; given that the particle content of scenario $\mathbf{B_{I}}$ is the same as the class-A scenarios, the new physics energy scale related to $\mathcal{O}_{5}^{(1)}$ and $\mathcal{O}_{5}^{(2)}$ for $\mathbf{B_{I}}$ can be read in the discussion above. In this scenario, if all the operators are generated with $\mathcal{O}(1)$ Wilson coefficients, $\Lambda$ should be the largest, i.e., $10^{12}$ GeV.
@@ 3.4 Naturally-induced small VEVs
  • In the $\mathbf{B_{I}}$ scenario, where we have two different quadruplets, there are both quartic linear terms ($\lambda_{5}$ and $\lambda_{6}$ terms in Table 3 for $\Phi_{1}$ and $\Phi_{2}$) and a mixed term proportional to $\lambda_{1}$. Again, in this case both VEVs can be induced by the two quartic terms, see Eq. (35), and are naturally suppressed. Another possibility is to induce one of the two VEVs with the the first two terms and the second one with the mixed term. As already mentioned, it is expected that the VEV induced by the last term, Eq. (36), is subdominant.
@@ 4.1.1 Production
Figure 5: Associated and pair production of multi-charged scalars present in our scenarios at the $\sqrt{13}$ TeV LHC.
We implement the models in SARAH Staub:2013tta and numerically compute the cross-sections in MadGraph5_amC@NLO Alwall:2014hca using the NNPDF23_nlo_as_0119 PDF set Ball:2013hta . The production cross sections for the charged scalars belonging to different multiplets at the LHC (for $\sqrt{s}=13$ TeV), as a function of the scalar mass, are shown in Fig. 5.
@@ 4.1.2 Decays
In the left (right) panel of Fig. 6 we show the branching ratio into different modes (proper decay length) of the doubly-charged scalar belonging to the multiplet $\mathbf{4}_{-1/2}$ of $\mathbf{A_{I}}$ (for all scenarios) as a function of the VEV of the multiplet. We take $M_{\Phi}=500$ and mass splitting $\Delta M=1\,(0,2.5)$ GeV in the left (right) panels. The grey curves on the left panel denote the cascade decays into pions, leptons and quarks.
@@
$v_{{\Phi^{\pm\pm}}}^{c}\simeq 65{\rm~{}KeV}\,\left(\frac{\kappa}{S_{2W^{\pm}}}\right)^{1/2}\,\left(\frac{\sum_{i}m_{i}^{2}}{0.05^{2}{\rm~{}eV}^{2}}\right)^{1/4}\,\left(\frac{500}{M_{\Phi^{\pm\pm}}}\right)^{1/2}\,.$ (40)
@@
Next, we focus on the decays of the triply-charged scalars. Depending on the mass splitting among the components, the dominant channels for their decays would be three-body states such as $llW$ or $WWW$, and two-body decays into ${\Phi^{\pm\pm}}W^{\pm\ast}$ and ${\Phi^{\pm\pm}}\pi^{\pm}$ for $M_{\Phi^{\pm\pm\pm}}>M_{\Phi^{\pm\pm}}$, which are similar to the decays of doubly-charged scalars into $\Phi^{+}\pi^{\pm}$ (one can obtain the expressions using appropriate scaling factors). The decay widths of the triply-charged scalar present in $\mathbf{4}_{-3/2}$ of $\mathbf{A_{II}}$ for $M_{\Phi^{\pm\pm\pm}}\gg M_{W}$ are
@@
$\Gamma({\Phi^{\pm\pm\pm}}\rightarrow W^{\pm}l^{\pm}l^{\pm})$ $=\frac{g^{2}}{6144\pi^{3}}\frac{M_{\Phi^{\pm\pm\pm}}\,\sum_{i}m_{i}^{2}}{v_{\Phi}^{2}}\,.$ (41)
The expressions for triply-charged scalars belonging to other multiplets can be obtained by rescaling the above expressions using the factors for couplings ${\Phi^{\pm\pm\pm}}\Phi^{\mp\mp}W^{\mp}$, ${\Phi^{\pm\pm}}W^{\mp}W^{\mp}$ and ${\Phi^{\pm\pm}}l^{\mp}l^{\mp}$, see Table 6.
@@
$v_{\phi^{\pm\pm\pm}}^{c}\simeq 18.5{\rm~{}KeV}\,\left(\frac{\kappa}{S_{2W^{\pm}}}\right)^{1/2}\,\left(\frac{\sum_{i}m_{i}^{2}}{0.05^{2}{\rm~{}eV}^{2}}\right)^{1/4}\,\left(\frac{500}{M_{\Phi^{\pm\pm}}}\right)\,.$ (42)
In the left (right) panel of Fig. 7 we show the branching ratio into different modes (proper decay length) of the triply-charged scalar of scenario $\mathbf{A_{II}}$ ($\mathbf{A_{II}},\mathbf{B_{I,II,III,IV}}$) as a function of the VEV of the multiplet. Being three-body decays, these decay lengths are much larger than those of doubly-charged scalars $\sim\mathcal{O}(0.1-1{\rm~{}mm})$ and lead to displaced vertices.
@@
$\Gamma_{\rm tot}({\Phi^{\pm\pm\pm\pm}})\sim\Gamma_{\rm tot}({\Phi^{\pm\pm\pm}})\,\frac{f(3)}{f(4)}\frac{g^{2}M_{{\Phi^{\pm\pm\pm\pm}}}^{2}}{M_{W}^{2}}\simeq 0.017\,\left(\frac{M_{{\Phi^{\pm\pm\pm\pm}}}}{500}\right)^{2}\,\Gamma_{\rm tot}({\Phi^{\pm\pm\pm}})\,,$ (43)

### v2

@@ 3.2.1 The $\mathbf{A_{I}}$ scenario
As expected, we obtain a pure massless would-be Nambu-Goldstone boson $z$, associated with the spontaneous breaking of the hypercharge, and a massive pseudo-Nambu-Goldstone boson $J$, related to the explicit breaking of the lepton number in one unit, that we identify as the pseudo-Majoron. Notice that its mass depends on the LNV quartic couplings of the potential and, as already mentioned, in the limit $(\lambda_{6},\lambda_{7},\lambda_{8})\rightarrow 0$, it becomes massless. The mixing angle $\alpha$ is given by
@@ 3.2.2 The $\mathbf{B_{I}}$ scenario
We now study the case of the $\mathbf{B_{I}}$ scenario as an example. We choose this specific class-B scenario since it is a natural extension of the $\mathbf{A_{I}}$ scenario discuss above. Indeed, in $\mathbf{B_{I}}$ we add to $\mathbf{A_{I}}$ a second multiplet, $\Phi_{2}=\mathbf{4}_{-3/2}$. Moreover, the study of such a scenario is particularly interesting since we can infer the results on the Majoron of the scenario $\mathbf{A_{II}}$ setting to zero all the scalar potential couplings in which $\Phi_{1}$ appears. In this scenario, as in all class-B scenarios, there are three $U(1)$ symmetries: $U(1)_{Y}$ of hypercharge, $U(1)_{L}$ of lepton number, and an additional $U(1)_{X}$ accidental symmetry. In the limit in which all the violating terms in the potential go to zero, we expect to have three massless Nambu-Goldstone bosons; if instead we turn on $\lambda_{1}$ (Eq. (16)) we get a massive Nambu-Goldstone boson i.e., the pseudo-Majoron $J$, associated with the explicit breaking of the lepton number. If $\lambda_{i}$ for $i=2,...,9$ are also different from zero (see Table 3), a second massive pseudo-Nambu-Goldstone boson $\Omega$ is obtained, related to the $U(1)_{X}$ explicit accidental symmetry breaking. As in Eq. (18), we can split the neutral scalar fields into their real and imaginary parts:
@@
As expected, we obtain a pure massless would-be Nambu-Goldstone boson $z$, associated with the spontaneous breaking of hypercharge, and two massive pseudo-Nambu-Goldstone bosons associated with the explicit breaking of the $U(1)_{L}$ and $U(1)_{X}$ symmetries, respectively. Even if at the leading order there is no dependence on the LNV potential coupling $\lambda_{1}$, in the limit of $U(1)_{X}$ conservation, namely when $\lambda_{i}$ with $i=2,...,9$ are zero, the pseudo-Majoron $J$ remains massive with $m_{J}^{2}\simeq\lambda_{1}v^{2}(v_{1}^{2}+v_{2}^{2})/(2\,v_{1}\,v_{2}\sqrt{3})$. Defining $(\tan\alpha)_{uv}$ as the tangent of mixing between the CP-odd scalars $u$ and $v$, we obtain:
@@
It is interesting to notice that whenever a massive pseudoscalar is included in our scenario, we might set lower bounds on the LNV couplings. Indeed, such particles cannot have a mass which is below 45 GeV, otherwise we would have observed them in $Z$-boson decays CMS:2022ett . If we have terms proportional to $\lambda{{}^{\prime\prime}}\Phi_{i}H^{3}$ in the potential ($\mathbf{A_{I}}$, $\mathbf{A_{II}}$, $\mathbf{B_{I}}$ scenarios), neglecting small corrections proportional to $v_{\Phi}$, we have
@@
This means that the lower limit on $\lambda^{\prime\prime}$, in the case of $v_{\Phi}\sim\mathcal{O}(1\,\,\rm{GeV})$ is $\lambda^{\prime\prime}>10^{-4}$. However, this limit can be further lowered taking a smaller $v_{i}$. In other scenarios, where we can have terms like $\lambda_{1}\Phi_{1}\Phi_{2}H^{2}$ and $\lambda^{\prime}\Phi_{i}^{2}H^{2}\,(i=1,2)$, there is no enhancement from $v_{i}$ in the denominator, and the requirement $m_{J}>45$ GeV gives $\lambda_{1},\lambda^{\prime}>10^{-1}$.
@@ 3.3 The VEVs of new scalars and the $\rho$ parameter
Table 4: 95% CL upper limits on the VEVs of the new scalar multiplets from the $\rho$ parameter and the possibility of inducing small VEVs in class-A scenarios.
First, let us consider the scenarios belonging to class-A, namely the ones in which we add a single scalar multiplet: $\Phi_{1}=\mathbf{4}_{-1/2}$ for $\mathbf{A_{I}}$ and $\Phi_{1}=\mathbf{4}_{-3/2}$ for $\mathbf{A_{II}}$. Recalling the general formula, Eq. (27), we get
@@
Using the experimental bound, we obtain the limits on the new VEVs, which are listed in Table 4. As expected, the BSM VEVs are always much smaller than the SM Higgs one. Using this result for which $v_{1}\sim\mathcal{O}(1\,\mathrm{GeV})$, it is also possible to estimate the energy scale of the UV complete theory, since from the Weinberg-like operators the neutrino mass can be written as $m_{\nu}\sim vv_{1}/\Lambda$ and $m_{\nu}\sim v_{1}^{2}/\Lambda$ (if the dimension-5 operators $\mathcal{O}_{5}^{(1)}$ and $\mathcal{O}_{5}^{(2)}$ are generated). Imposing $m_{\nu}\gtrsim\mathcal{O}(0.05\,\mathrm{eV})$, we can estimate the order of magnitude of the energy scale $\Lambda$. Since we do not know in principle which of the operators gives the leading contribution to neutrino masses, we show here the mass scales $\Lambda^{(i)}$ one would obtain if only one of the operators exists. Taking the Wilson coefficients to be $\mathcal{O}(1)$, we obtain for perturbative Wilson coefficients $\Lambda^{(1)}\lesssim 10^{12}$ GeV for $\mathcal{O}_{5}^{(1)}$, which is present for both scenarios $\mathbf{A_{I,II}}$ and $\Lambda^{(2)}\lesssim 10^{10}$ GeV for $\mathcal{O}_{5}^{(2)}$, which is present for the scenario $\mathbf{A_{I}}$ but not for the scenario $\mathbf{A_{II}}$.
@@
Figure 4: Allowed regions for the new scalar multiplet VEVs in class-$\mathbf{B}$ scenarios computed using the 95% CL limits on the $\rho$ parameter, $\rho=1.00017\pm 0.00025$ ParticleDataGroup:2022pth . Two different structures emerge: hyperbolae for $\mathbf{B_{I,III,V,VI}}$ and ellipses for $\mathbf{B_{II,IV}}$, see text for details.
@@
When the allowed region is an ellipse (scenarios $\mathbf{B_{II}}$ and $\mathbf{B_{IV}}$), the values of the two semi-axes $a_{1,2}$ correspond to the maximum values that the VEVs can assume. On the other hand, when the allowed region is the space between two hyperbolae (scenarios $\mathbf{B_{I}}$, $\mathbf{B_{III}}$, $\mathbf{B_{V}}$ and $\mathbf{B_{VI}}$), the BSM VEVs can in principle also have large values if there is some partial cancellation among the contributions. However, they cannot exceed the limit imposed by the Fermi constant, namely 174 GeV. Moreover, given that the mass of the top quark is proportional to the SM Higgs VEV $v$, by perturbativity the latter is not allowed to be too small. Thus, the new VEVs $v_{1}$ and $v_{2}$ in the $\mathbf{B_{I,III,V,VI}}$ scenarios also cannot exceed roughly $\mathcal{O}(10\,\text{GeV})$.
Analogous to the case of class-A scenarios, we can deduce the energy scale of the new physics by imposing that neutrino masses are $m_{\nu}\gtrsim\mathcal{O}(0.05\,\mathrm{eV})$ and that the Wilson coefficients are of $\mathcal{O}(1)$. Since for class-B scenarios we can have the dimension-5 operator $\mathcal{O}_{5}^{(3)}$ from which $m_{\nu}\sim v_{1}v_{2}/\Lambda$, we obtain for all the scenarios, an energy scale $\Lambda\lesssim\mathcal{O}(10^{10})$ GeV. The semi-axes of the conics which represent the allowed region for the new scalar VEVs obtained from EWPT for the class-B scenarios are listed in Table 5. Notice that in scenario $\mathbf{B_{I}}$, the Wilson coefficients for $\mathcal{O}_{5}^{(1)}$ and $\mathcal{O}_{5}^{(2)}$ related to neutrino masses are not zero; given that the particle content of scenario $\mathbf{B_{I}}$ is the same as the class-A scenarios, the new physics energy scale related to $\mathcal{O}_{5}^{(1)}$ and $\mathcal{O}_{5}^{(2)}$ for $\mathbf{B_{I}}$ can be read in the discussion above. In this scenario, if all the operators are generated with $\mathcal{O}(1)$ Wilson coefficients, $\Lambda$ should be the largest, i.e., $10^{12}$ GeV.
@@ 3.4 Naturally-induced small VEVs
  • In the $\mathbf{B_{I}}$ scenario, where we have two different quadruplets, there are both quartic linear terms ($\lambda_{5}$ and $\lambda_{6}$ terms in Table 3 for $\Phi_{1}$ and $\Phi_{2}$) and a mixed term proportional to $\lambda_{1}$. Again, in this case both VEVs can be induced by the two quartic terms, see Eq. (33), and are naturally suppressed. Another possibility is to induce one of the two VEVs with the first two terms and the second one with the mixed term. As already mentioned, it is expected that the VEV induced by the last term, Eq. (34), is subdominant.
@@ 4.1.1 Production
Figure 5: Associated (dashed curves) and pair production (solid curves) of multi-charged scalars present in our scenarios at the $\sqrt{13}$ TeV LHC.
We implement the models in SARAH Staub:2013tta and numerically compute the cross-sections in MadGraph5_amC@NLO Alwall:2014hca using the NNPDF23_nlo_as_0119 PDF set Ball:2013hta . The production cross sections for the charged scalars belonging to different multiplets at the LHC (for $\sqrt{s}=13$ TeV), as a function of the scalar mass, are shown in Fig. 5. It can be seen that the pair production cross section for the scalars is directly proportional to the charge of the scalars, therefore, in all cases, we see that $\sigma(pp\rightarrow\phi^{n\pm}\phi^{n\mp})>\sigma(pp\rightarrow\phi^{m\pm}\phi^{m\mp})$ for $n>m\geq 2$, where $n,m$ indicate the scalar charges. However, this is not true for the associated production that takes place via $W^{\pm}$ exchange. From Table 6, one can see that the numerical factors associated with $\phi^{\pm\pm}\phi^{\mp}W^{\mp}$ couplings for different scalar multiplets are greater or equal to those of $\phi^{\pm\pm\pm}\phi^{\mp\mp}W^{\mp}$, therefore we obtain $\sigma(pp\rightarrow\phi^{n\pm}\phi^{(n-1)\mp})\geq\sigma(pp\rightarrow\phi^{m\pm}\phi^{(m-1)\mp})$ for $n<m$ with ${n,m\geq 2}$.
@@ 4.1.2 Decays
In the left (right) panel of Fig. 6 we show the branching ratio into different modes (proper decay length) of the doubly-charged scalar belonging to the multiplet $\mathbf{4}_{-1/2}$ of $\mathbf{A_{I}}$ (for all scenarios) as a function of the VEV of the multiplet. We take $M_{\Phi}=500$ GeV and mass splitting $\Delta M=1\,(0,2.5)$ GeV in the left (right) panels. The grey curves on the left panel denote the cascade decays into pions, leptons and quarks.
@@
$v_{{\Phi^{\pm\pm}}}^{c}\simeq 65{\rm~{}KeV}\,\left(\frac{\kappa}{S_{2W^{\pm}}}\right)^{1/2}\,\left(\frac{\sum_{i}m_{i}^{2}}{0.05^{2}{\rm~{}eV}^{2}}\right)^{1/4}\,\left(\frac{500{\rm~{}GeV}}{M_{\Phi^{\pm\pm}}}\right)^{1/2}\,.$ (38)
@@
Next, we focus on the decays of the triply-charged scalars. Depending on the mass splitting among the components, the dominant channels for their decays would be three-body states such as $llW$ or $WWW$, and two-body decays into ${\Phi^{\pm\pm}}W^{\pm\ast}$ and ${\Phi^{\pm\pm}}\pi^{\pm}$ for $M_{\Phi^{\pm\pm\pm}}>M_{\Phi^{\pm\pm}}$, which are similar to the decays of doubly-charged scalars into $\Phi^{+}\pi^{\pm}$ (one can obtain the expressions using appropriate scaling factors). The decay widths of the triply-charged scalar present in $\mathbf{4}_{-3/2}$ of $\mathbf{A_{II}}$ for $M_{\Phi^{\pm\pm\pm}}\gg M_{W}$ are Arbelaez:2019cmj ; Pan:2019wwv
@@
$\Gamma({\Phi^{\pm\pm\pm}}\rightarrow W^{\pm}l^{\pm}l^{\pm})$ $=\frac{g^{2}}{3072\pi^{3}}\frac{M_{\Phi^{\pm\pm\pm}}^{3}\,\sum_{i}m_{i}^{2}}{v_{\Phi}^{2}\,M_{W}^{2}}\,.$ (39)
Note that for large scalar masses, $M_{\Phi^{\pm\pm\pm}}\gg m_{W}$, the decay should be present in the gauge-less limit (Goldstone Boson Equivalence Theorem). [^7] The expression for $\Gamma({\Phi^{\pm\pm\pm}}\rightarrow W^{\pm}l^{\pm}l^{\pm})$ in Eq. (4.1.2) differs by a factor of $2(M_{\Phi^{\pm\pm\pm}}^{2}/M_{W}^{2})$ from that in Refs. Babu:2009aq ; Ghosh:2017jbw , where the decay rate vanishes in the gauge-less limit. The expressions for triply-charged scalars belonging to other multiplets can be obtained by rescaling the above expressions using the factors for couplings ${\Phi^{\pm\pm\pm}}\Phi^{\mp\mp}W^{\mp}$, ${\Phi^{\pm\pm}}W^{\mp}W^{\mp}$ and ${\Phi^{\pm\pm}}l^{\mp}l^{\mp}$, see Table 6.
@@
$v_{\phi^{\pm\pm\pm}}^{c}\simeq 55.4{\rm~{}KeV}\,\left(\frac{\kappa}{S_{2W^{\pm}}}\right)^{1/2}\,\left(\frac{\sum_{i}m_{i}^{2}}{0.05^{2}{\rm~{}eV}^{2}}\right)^{1/4}\,\left(\frac{500{\rm~{}GeV}}{M_{\Phi^{\pm\pm}}}\right)^{1/2}\,.$ (40)
In the left (right) panel of Fig. 7 we show the branching ratio into different modes (proper decay length) of the triply-charged scalar of scenario $\mathbf{A_{II}}$ ($\mathbf{A_{II}},\mathbf{B_{I,II,III,IV}}$) as a function of the VEV of the multiplet. Being three-body decays, these decay lengths $\sim\mathcal{O}(0.01-1{\rm~{}mm})$ are much larger than those of doubly-charged scalars, and may lead to displaced vertices.
@@
$\Gamma_{\rm tot}({\Phi^{\pm\pm\pm\pm}})\sim\Gamma_{\rm tot}({\Phi^{\pm\pm\pm}})\,\frac{f(3)}{f(4)}\frac{g^{2}M_{{\Phi^{\pm\pm\pm\pm}}}^{2}}{M_{W}^{2}}\simeq 0.017\,\left(\frac{M_{{\Phi^{\pm\pm\pm\pm}}}}{500{\rm~{}GeV}}\right)^{2}\,\Gamma_{\rm tot}({\Phi^{\pm\pm\pm}})\,,$ (41)